Consider a mass-spring system. The spring has a spring constant of 2.27e+3 N/m. On the end of the spring is a mass of 2.20 kg. At the moment that the block is at the system's equilibrium position, it has velocity of 1.89 m/s. What will be the maximum displacement of the block from equilibrium if we ignore all friction.
Answer #1

Energy conservation =

Spring potential energy equal to the kinetic energy

1/2kA^2 = 1/2m*V^2

A= V√(m/k) = 1.89*√(2.2/2.27*10^3)

A= 0.0588m Or 5.89 cm